# Scaling Large Language Model Training on Frontier with Low-Bandwidth Partitioning

### v1

@@ I Abstract
Scaling up Large Language Model(LLM) training involves fitting a tremendous amount of training parameters across a limited number of workers. However, methods like ZeRO-3 that drastically reduce GPU memory pressure often incur heavy communication to ensure global synchronization and consistency. Established efforts such as ZeRO++ use secondary partitions to avoid inter-node communications, given that intra-node GPU-GPU transfer generally has more bandwidth and lower latency than inter-node connections. However, as more capable infrastructure like Frontier, equipped with AMD GPUs, emerged with impressive computing capability, there is a need for investigations on the hardware topology and to develop targeted strategies to improve training efficiency. In this work, we propose a collection of communication and optimization strategies for ZeRO++ to reduce communication costs and improve memory utilization. In this paper, we propose a 3-level hierarchical partitioning specifically for the current Top-1 supercomputing cluster, Frontier, which aims at leveraging various bandwidths across layers of communications (GCD-GCD, GPU-GPU, and inter-node) to reduce communication overhead. For a 20B GPT model, we observe a 1.71x increase in TFLOPS per GPU when compared with ZeRO++ up to 384 GCDs and a scaling efficiency of 0.94 for up to 384 GCDs. To the best of our knowledge, our work is also the first effort to efficiently optimize LLM workloads on Frontier AMD GPUs.  [^1] This research is supported in part by NSF grants #1818253, #1854828, #2018627, #2311830, #2312927, #2415201, and XRAC grant #NCR-130002.  [^2] This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DEAC05-00OR22725.
@@ II-A Motivation
Recently, we have also witnessed growing attention paid to the AMD compute stack. For example, Frontier [16], the Top-1 supercomputing cluster, is equipped with compute nodes that have four MI250X GPUs, connected using Infinity Fabric within a node and Slingshot 11 [17] across nodes. Communications are conducted through RCCL. System topology is detailed in Section IV. Given the low-bandwidth configuration compared to DGX systems, LLM training on such platforms has been an under-studied area, which leaves room for investigation and improvements.
@@ II-D Contributions
    To the best of our knowledge, this is the first work that adapts ZeRO++ and validates its effectiveness on AMD GPUs. Also, this is the first work that proposed improved hierarchical partitioning on top of ZeRO++, which features a software-hardware co-design on the Top-1 supercomputing infrastructure.
@@ IV System Architecture Analysis
This section presents a comprehensive system architecture analysis of popular large-scale HPC clusters used for training large language models (LLMs). Our comparison focuses on clusters built around different GPU vendors, primarily NVIDIA and AMD. Additionally, we analyze the compute node topology of the current TOP-1 Frontier supercomputing cluster hosted by Oak Ridge National Laboratory. We begin by examining a DGX-A100 node designed for NVIDIA GPUs.
@@
Next, we examine Frontier, the current top-ranked HPE Cray EX supercomputing cluster hosted by Oak Ridge National Laboratory. Each compute node in Frontier contains 4 AMD MI250X GPUs, each equipped with 2 Graphic Computing Dies (GCDs) and 128GB of HBM memory with a bandwidth of 1.6TB/s. Within each MI250X are four Infinity Fabric links, providing a total GCD-to-GCD bandwidth of 200GB/s. Each pair of MI250X GPUs is connected by two Infinity Fabric links (100GB/s) for adjacent pairs and one link (50GB/s) for cross pairs. Inter-node connections are established using 4 HPE Slingshot ports, delivering a total bandwidth of 100GB/s.
@@ IX Conclusion
Recent research efforts to decrease communication overhead in large-scale LLM training have primarily focused on the bandwidth gap between intra-node and inter-node communication, often conducted on NVIDIA platforms. However, with the emergence of more capable infrastructure featuring AMD GPUs, a meticulous co-design of software optimization and underlying hardware topology is essential for achieving greater efficiency. In this work, we propose a dedicated 3-level topology-aware hierarchical partitioning strategy tailored for the Frontier supercomputing cluster, the current Top-1 system. This strategy distributes training parameters across different layers of devices to fully utilize the interconnect bandwidth between GCDs, GPUs, and nodes. We implemented this protocol and validated it across various models and scales, achieving up to a 139.8% increase over ZeRO-3 and 70.7% increase over ZeRO++ in TFLOPS per GPU for a 20B model on 384 GPUs.

### v2

@@ I Abstract
Scaling up Large Language Model(LLM) training involves fitting a tremendous amount of training parameters across a limited number of workers. However, methods like ZeRO-3 that drastically reduce GPU memory pressure often incur heavy communication to ensure global synchronization and consistency. Established efforts such as ZeRO++ use secondary partitions to avoid inter-node communications, given that intra-node GPU-GPU transfer generally has more bandwidth and lower latency than inter-node connections. However, as more capable infrastructure like Frontier, equipped with AMD GPUs, emerged with impressive computing capability, there is a need for investigations on the hardware topology and to develop targeted strategies to improve training efficiency. In this work, we propose a collection of communication and optimization strategies for ZeRO++ to reduce communication costs and improve memory utilization. In this paper, we propose a 3-level hierarchical partitioning specifically for the current 2nd ranked supercomputing cluster, Frontier, which aims at leveraging various bandwidths across layers of communications (GCD-GCD, GPU-GPU, and inter-node) to reduce communication overhead. For a 20B GPT model, we observe a 1.71x increase in TFLOPS per GPU when compared with ZeRO++ up to 384 GCDs and a scaling efficiency of 0.94 for up to 384 GCDs.  [^1] This research is supported in part by NSF grants #1818253, #1854828, #2018627, #2311830, #2312927, #2415201, and XRAC grant #NCR-130002.  [^2] This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DEAC05-00OR22725.
@@ II-A Motivation
Recently, we have also witnessed growing attention paid to the AMD compute stack. For example, Frontier [16], the current 2nd ranked supercomputing cluster, is equipped with compute nodes that have four MI250X GPUs, connected using Infinity Fabric within a node and Slingshot 11 [17] across nodes. Communications are conducted through RCCL. System topology is detailed in Section IV. Given the low-bandwidth configuration compared to DGX systems, LLM training on such platforms has been an under-studied area, which leaves room for investigation and improvements.
@@ II-D Contributions
    To the best of our knowledge, this is the first work that adapts ZeRO++ and validates its effectiveness on AMD GPUs. Also, this is the first work that proposed improved hierarchical partitioning on top of ZeRO++, which features a software-hardware co-design on the 2nd ranked supercomputing infrastructure.
@@ IV System Architecture Analysis
This section presents a comprehensive system architecture analysis of popular large-scale HPC clusters used for training large language models (LLMs). Our comparison focuses on clusters built around different GPU vendors, primarily NVIDIA and AMD. Additionally, we analyze the compute node topology of the current 2nd ranked Frontier supercomputing cluster hosted by Oak Ridge National Laboratory. We begin by examining a DGX-A100 node designed for NVIDIA GPUs.
@@
Next, we examine Frontier, the current 2nd ranked HPE Cray EX supercomputing cluster hosted by Oak Ridge National Laboratory. Each compute node in Frontier contains 4 AMD MI250X GPUs, each equipped with 2 Graphic Computing Dies (GCDs) and 128GB of HBM memory with a bandwidth of 1.6TB/s. Within each MI250X are four Infinity Fabric links, providing a total GCD-to-GCD bandwidth of 200GB/s. Each pair of MI250X GPUs is connected by two Infinity Fabric links (100GB/s) for adjacent pairs and one link (50GB/s) for cross pairs. Inter-node connections are established using 4 HPE Slingshot ports, delivering a total bandwidth of 100GB/s.
@@ VIII Related Work
There have also been several efforts on optimizing LLM workloads on Frontier. [31, 32, 33]
@@ IX Conclusion
Recent research efforts to decrease communication overhead in large-scale LLM training have primarily focused on the bandwidth gap between intra-node and inter-node communication, often conducted on NVIDIA platforms. However, with the emergence of more capable infrastructure featuring AMD GPUs, a meticulous co-design of software optimization and underlying hardware topology is essential for achieving greater efficiency. In this work, we propose a dedicated 3-level topology-aware hierarchical partitioning strategy tailored for the Frontier supercomputing cluster, the current 2nd ranked system. This strategy distributes training parameters across different layers of devices to fully utilize the interconnect bandwidth between GCDs, GPUs, and nodes. We implemented this protocol and validated it across various models and scales, achieving up to a 139.8% increase over ZeRO-3 and 70.7% increase over ZeRO++ in TFLOPS per GPU for a 20B model on 384 GPUs.